Solve the following equation for x. [Do not report answers in decimal – leave them in radical form, and be sure to simplify radicals.
2(x – 5)^2 = 3

Answer by ankor@dixie-net.com(22740)

Solve the following equation for x. [Do not report answers in decimal – leave them in radical form, and be sure to simplify radicals.
:
2(x – 5)^2 = 3
:
divide both sides by 2 & use the latter half of the "Complete the Square" method:
(x- 5)^2 = (3/2)
:
Find the square root of both sides
x - 5 = +/-SqRt(3/2)
:
x = 5 + SqRt(3/2)
and
x = 5 - SqRt(3/2)